Project

General

Profile

Meeting 3-6-2015

Date: 3rd June 2015
Time: 15:00 UK time
Attendance: Ed O, Chris Hunt Adam, Ray, Melissa, Paul
Apologies:

Agenda and Minutes

Actions from Last Meeting

  • Kalman doesn't hasn't been checked on straights and there are small effects to be tested. - Chris Hunt deviations from an ideal field will be looked at but will be postponed until after straights and small effects have been considered.
  • Real data unpacking broken. A proper summary of the issue, its status and timeline to be presented this week. - Paul Ed o.and MU - Data unpaker is now almost fixed. Readout data had fixes on Monday that seems to now work. One broken VLSB chip will not read out. Ed is fixing. VLSB master will be modified to get trigger reading into DAQ so that nu. of triggers can be checked between trackers and other detectors. This functionality will be added to data structure and will not effect VLSB data structure. This will be tested. Paul K
  • Software update from Paul this meeting - PK. - nothing new to say. Mainly unpacking progress.
  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick. - Ongoing
  • Add to John Govans' rota for daily walk round to test our setup. - Pierrick. - Ongoing
  • Look at docs to determine what to release as MICE notes and how - MU - ONGOING not highest priority
  • EO to speak to C. Heidt when ADC/VLPC and ADC calib and MAUS transition performing calibrations. - Scheduled for Friday - Happened breifly and work plan agreed.
  • CDB script to up/download is done, needs to add to the MAUS mapper to use the script as MUAS is run. - C. Heidt
  • Multiple tracks per Tracker. - Durga - no progress.
  • Add online recon to trunk - MU - this week deadline
  • Fix C&M code - Pierrick - Fixed!
  • Check with DL to move interlocks after first user run - MU -Met breifly will chat again.
  • Email update on Thursday of MDR3 result. - EO -DONE.

0. MDR3

Most of day spent getting things going in the right order and fixing bugs.
OM was working and data was seen.
Unpacking was worked on in parrallel.
Post Mortem recovery plans today and yesterday:
VLSB readout time saturates at 300 particle triggers per spill for half a tracker. Worrying as were looking at 150 particles per spill per tracker. Will look at speeding up later.
Final calib of cryo1 being done today so a bit more post mortem work required.
Trigger code being implemented.

1. Status of hardware, commissioning work and ongoing plan EO, AB, CMW, PR, MU

  • Cryo 2 waveguided.
  • Cryo 1 will be waveguided tomorrow.
  • Cryo 1 commissioning done.
  • By end of tomorrow calibration & waveguiding of US tracker will be done.
  • Over the next few days US tracker with data will be tested from Thursday night with TOF0 and align windows.
  • TOF1 was damaged as it was craned out. No idea how long to fix as yet and harder for us as we don't know what the
  • Run Plan:
    - Thurs timing window with TOF0
    - Fri commissioning US tracker with beam.
    - Sat and Sunday following on from Friday if more time is needed.
  • DS commissioning to go on in the days as US data is in the night.

2. Software: PK, AD, CH

In particular this week:
  • Data unpacking
  • Kalman
  • DAQ in data structure
  • CDB mapping and bias calibrations

As in minutes:

  • Chris Hunt deviations from an ideal field will be looked at but will be postponed until after straights and small effects have been considered.
  • Real data unpacking broken. A proper summary of the issue, its status and timeline to be presented this week. - Paul Ed o.and MU - Data unpaker is now almost fixed. Readout data had fixes on Monday that seems to now work. One broken VLSB chip will not read out. Ed is fixing. VLSB master will be modified to get trigger reading into DAQ so that nu. of triggers can be checked between trackers and other detectors. This functionality will be added to data structure and will not effect VLSB data structure. This will be tested. Paul K

Geometry, Calibrations, Configurations, CDB - C. Heidt

No update

Monte Carlo

No update

In a good state waiting on Durga for multiple tracks per Tracker.

Pattern Recognition - AD,

  • No method to get multiple tracks per trigger.

Online Displays - MU

  • 3/4 of the plots are now in and working and were used for mock data challenge. The other are the non-immediately-essentials and are coming soon.

Performance and Analysis -PK

  • Durga has been working on the mappers and getting rid of JASOn and is making good progress.
  • Adam has been also pulling JASON out and this is also going well and is speeding things up.
  • Profiling will be rerun after this work has been.
  • Currently MAUSwide online recon didnt work last night and no idea why as it was working at MDR3.

Issues

Tracker Software Paper - AD, PK DR

Documentation - PK, AD, C. Heidt

C. Heidt: MC documentation coming along

4. Optimising the DAQ and Controls and Monitoring - EO, AB, PR

  • 50% done, but not enough time to fully go over before beam. Davids original code for MDR3 and new code to come in after first data.

5. MICE hall, inc Services hardware and requirements : CM

  • see schedule in documents.

6. Step IV Analysis Readiness - MU

  • Alignment data run requests submitted.
  • Unpacking and speed are biggest problems (PK)

7. Risks

7.1 Hardware

  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Northside PRY (not us, but knock effect to us)

7.2 Software

  • Not yet able to test against multiple tracks per trigger with MC -Action for Durga
  • Real data unpacking broken DA

8. Tracker Documentation

We need to produce documentation for the control room and ISIS, for Steve B. We cannot necessarily get that far with operations and startup as we are not yet working with the CandM as we will but we can write something and finalise safety details.

8.1 Startup Procedure

8.2 Operations Procedure

8.3 Tracker Online Emergency Procedure

  • We have been asked to provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks

5. Other Items

  • Installation Manual - GB MU - ongoing
  • LED system - EO - External used for calib, Internal not used since Dec but will be used after gain and pedestals determined.
  • Hall probes - MU - Electronics readout devised and computing needs obtained. Cabling to be made.

AoB

New Actions

  • Kalman doesn't hasn't been checked on straights and there are small effects to be tested. Deviations from an ideal field will be looked at but will be postponed until after straights and small effects have been considered.. - Chris Hunt
  • Real data unpacking. A proper summary of the issue, its status and timeline to be presented this week. - Paul Ed O.and MU - Data unpaker is now almost fixed and VLSB master modification and One broken VLSB chip will not read out.
  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick. - Ongoing
  • Add to John Govans' rota for daily walk round to test our setup. - Pierrick. - Ongoing
  • Look at docs to determine what to release as MICE notes and how - MU - ONGOING not highest priority
  • CDB script to up/download is done, needs to add to the MAUS mapper to use the script as MUAS is run. - C. Heidt
  • Multiple tracks per Tracker. - Durga -
  • Add online recon to trunk - MU
  • Check with DL to move interlocks after first user run - MU

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English